OpenStudy (anonymous):

simplify (√3) (^5√3)

OpenStudy (anonymous):

HELP

OpenStudy (misty1212):

HI!!!

OpenStudy (anonymous):

HAI

OpenStudy (misty1212):

is it \[\huge\sqrt{3}\sqrt[5]{3}\]

OpenStudy (anonymous):

Yea

OpenStudy (misty1212):

you can think of it as \[\huge x^{\frac{1}{2}}\times x^{\frac{1}{5}}\] then add the exponents

OpenStudy (anonymous):

ooohhhh ok :) thanks i got this

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]